# AIGB: Generative Auto-bidding via Diffusion Modeling

### v2

@@ 3. AIGB PARADIGM for AUTO-BIDDING
To thoroughly investigate the distribution of impression opportunities, we have carried out a series of statistical analyses. As depicted in Figure 3(a), a representative example, comprising three time steps sampled from the online advertising system, shows that CE exhibits a power-law decline. This decline becomes evident as the number of winning impressions increases at each time step, with decay rates varying temporally due to the impressions’ heterogeneous nature. Consequently, bidding agents adjust their bids to align with these distributions, targeting impressions that exceed the CE ratio and the bidding strategy effectively boils down to selecting a specific number of winning impressions per time step. Additionally, Figure 3(b) demonstrates that the average impression cost remains relatively stable, fluctuating by less than 5%. This stability allows us to approximate the total cost $c_{t}$ at each time step using the average cost $\bar{c}=\frac{1}{N}\sum_{i=0}^{N}c_{i}$, where $N$ represents the number of winning opportunities of the whole episode. Consequently, we can express $c_{t}=n_{t}\cdot\bar{c}$, where $n_{t}$ indicates the count of winning impressions at time step $t$.
This analysis leads us to propose a trajectory-wise perspective for auto-biding, in which devising an effective bidding strategy is essentially equivalent to identifying a proper state trajectory that accommodates $c_{t}$ and integrates other performance-related statuses alongside. This insight enables us to tackle the auto-bidding problem using a hierarchical paradigm, which prioritizes the state sequence optimization and subsequently generates actions that are consistent with this optimized trajectory.
The hierarchical paradigm divides auto-bidding into two supervised learning problems, offering a suite of advantages that include enhanced interpretability and increased stability during the training process. For state sequence optimization, our method directly models the joint distribution of the entire bidding trajectory alongside its associated returns and generates the trajectory distribution conditioned on the target return. This approach enables us to address the key auto-bidding challenges by employing SOTA generative algorithms. In this paper, we present an implementation that utilizes Denoising Diffusion Probabilistic Models (DDPM).
(a) Cost Effectiveness Curve Samples from Three Steps
(b) Impression Cost Fluctuation at Different Time Steps
Figure 3. Statistical Results From Online Advertising System.
@@ A.5. Analytical Results for Action Control
A.6. Theoretical Analysis

### v4

@@ 3. AIGB PARADIGM for AUTO-BIDDING
To thoroughly investigate the auto-bidding problem, we conducted a series of statistical analyses of bidding trajectories, with detailed information available in appendix LABEL:statistical_analysis. These analyses provide us with the insight that devising an effective bidding strategy is essentially equivalent to optimizing a state trajectory. Armed with this insight, we propose a hierarchical paradigm for auto-bidding that prioritizes the state trajectory optimization and subsequently generates actions aligned with the optimized trajectory.
For state trajectory optimization, we can employ a generative model to capture the joint distribution of the entire bidding trajectory and its associated returns, subsequently generating the trajectory distribution conditioned on the desired return. This approach enables us to address key auto-bidding challenges by employing SOTA generative algorithms. This paper presents an implementation that utilizes Denoising Diffusion Probabilistic Models (DDPM). For action generation, several off-the-shelf methods can be utilized to predict the proper action given the target state trajectory. In this paper, we apply a widely used inverse dynamics model. The hierarchical paradigm divides auto-bidding into two supervised learning problems, offering several advantages that include enhanced interpretability and increased stability during the training process.
@@ A.5. Analytical Results for Action Control
A.6. Statistical Analyses for Bidding Trajectory
The study by (Hao et al., 2020) indicates that CE follows a power-law decline as the number of winning impressions increases. Our statistical analysis confirms that this finding holds true at every discrete time step, with decay rates varying temporally due to the heterogeneous nature of the impressions. Figure 8(a) shows three steps sampled from the online advertising system, From which another key insight is that the optimal bidding strategy’s $ce*$ is equivalent to selecting a specific number of winning impressions per time step. We denote the number at time step $t$ as $n_{t}$.
Another finding illustrated in Figure 8(b) is that the costs of impressions remain relatively stable throughout the total episode, fluctuating by less than 5%. This stability allows us to approximate the cost of each impression $c_{i}$ with the average cost $\bar{c}=\frac{1}{N}\sum_{i=0}^{N}c_{i}$, where $N$ represents the number of winning impressions of the total episode. Therefore, the total cost at each time step $c_{t}=n_{t}\cdot\bar{c}$. In auto-bidding modeling, $c_{t}$ can be calculated from the state trajectory by using the difference in the remaining budget between two consecutive steps. Consequently, we can conclude that the optimal strategy correlates to a specific state trajectory.
A.7. Theoretical Analysis
@@
(a) Cost Effectiveness Curve Samples from Three Steps
(b) Impression Cost Fluctuation at Different Time Steps
Figure 8. Statistical Results From Online Advertising System.